OpenStudy (anonymous):

What is wrong with the equation E =M(c)squared? I heard that as mass approaches the speed of light, its inertia approaches infinity. Yet no matter what M I put in with c being the speed of light comes out with E being infinity. What am I missing?

OpenStudy (theeric):

I could be wrong, but I think that \(E\) is the internal energy associated with the mass. Are you saying that your calculator says that \(E=\infty\)? That might be a calculator issue. Since \(c\approx 3.00\times10^8\rm\ m/s\), try using \(c=3\) and just multiply your answer by \(10^8\).

OpenStudy (irishboy123):

\[E_{0} = m_{0}c^{2}\] This addresses the body's rest mass (Mo) and rest energy (Eo) and is the simplest expression of the mass-energy equivalence. c is in this way just another one of those fab numbers like π or e. in terms of your query, if the body moves something else happens: \[m _{v} = \frac{ m _{0} }{\sqrt{\left( 1 - \frac{ v^{2} }{ c^{2} } \right)} }\] as v, the velocity of the object, tends towards c, v/c tends to 1 so the denominator tends to zero, meaning Mv tends to infinity. this explains why matter cannot exceed the speed of light whereas massless photons can ping around at c. the extra energy you put in trying to accelerate the body also increases its mass as mass is energy. Combine the 2 and you get this: \[E = \frac{ m_{0}c^{2} }{\sqrt{\left( 1 - \frac{ v^{2} }{ c^{2} } \right)} }\] Same idea - as the object speeds up, both its innate energy content and its mass increase.

OpenStudy (anonymous):

Correct.. E = mc^2 is only when an object is AT REST.. it basically says that mass can be converted to energy and vice versa.. on the other hand, special relativity also postulates that nothing can go faster than speed of light.. if that is true.. what would happen if u kept pumping more and more energy into a body? cause the kinetic energy should increase right? that means speed should increase? well , in reality both mass and speed increases.. the mass increase is insignificant at low speeds, but then becomes dominant, and tends to become infinity as velocity tends to speed of light

OpenStudy (anonymous):

Correct.. E = mc^2 is only when an object is AT REST.. ????? why do you think so?

OpenStudy (anonymous):

cause i know so.. that equation uses REST mass of an object.. :P

OpenStudy (anonymous):

if the mass is moving, then along with that, there should be kinetic energy added to the right side,.

OpenStudy (anonymous):

but E=mc^2 is equal to kinetic energy plus its rest energy.. it never applies at rest.

OpenStudy (anonymous):

nope.. mc^2.. is the energy associated with the MASSS of the particle .. so if the particle is moving, it has additional kinetic energy!

OpenStudy (anonymous):

it is associated with RELATIVE mass but not rest mass? do you know what the relative mass is?

OpenStudy (anonymous):

I really don't know what you are trying to conclude???

OpenStudy (anonymous):

Thank you Saeeddiscover and Mashy. Your answers helped me to see MY error. When the velocity is equal to 0, then E=M. When the velocity is greater than 0, then E can represent the rest energy and the kinetic energy. And E is infinity when c is the speed of light. I love science but the understanding doesn't come easy to me.